# One for All: Towards Training One Graph Model for All Classification Tasks

### v2

@@ B.1.2 PubMed
Cora is a citation network contains papers and their citation relationship in biomedical domain. The raw text data of the PubMed dataset was collected from the GitHub repository provided in Chen et al. (2023). All nodes and edges are similar to Cora dataset and we use the exact the same processing procedure as Cora dataset. Because its original literature categories are diabetes, experimental/diabetes, type 1/diabetes, type 2, which are overly simple and very difficult for our BERT-based LLM to distinguish, we asked ChatGPT to generate a detailed description of each category.
@@
Cora is a citation network contains papers and their citation relationship collected from arxiv platform. The raw text data of the ogbn-arxiv was collected using the same protocol as the GitHub repository provided in Prodigy (Huang et al., 2023). For ogbn-arxiv, we only evaluate the model on the node classification task. All nodes and edges are similar to Cora dataset and we use the exact the same processing procedure as Cora dataset except that the description for each category is directly obtained from the Prodigy.
@@ B.1.4 Wiki-CS
Wiki-CS is a Internet link network with each node represent a Wikipedia page and each edge represent the reference link. The raw text of Wiki-CS dataset was collected from the official website (Mernyei & Cangea, 2020). The raw text feature of a node is the name and content of an entry in Wikipedia. Each node’s label corresponds to the category of the entry. We evaluate Wiki-CS on node calssification task. In table 9, we show a processed example on Wiki-CS dataset using the processing protocol proposed in OFA.
@@
Node type
Text feature
Input graph node
feature node. wikipedia entry name: $<$entry name$>$. Entry content: $<$entry content$>$
Input graph edge
feature edge. wikipedia page link.
@@
Class node
prompt node. wikipedia entry category: $<$category name$>$
Prompt node
prompt node. node classification of wikipedia entry category.
@@ B.1.5 FB15K237
FB15K237 is a kowledge graph that contains knowledge base relation triples and textual mentions of Freebase entity pairs. The raw text data of nodes in FB15K237 was collected from GitHub repository [^1] https://github.com/villmow/datasets_knowledge_embedding/tree/master. The raw text feature of a node is the name of the relation entity and its description. The raw text feature of an edge is the type of the relation between two entities. We provide an example of processed data example in Table 10.
@@
Node type
Text feature
Input graph node
feature node. entity and entity description: <entity name>. <entity description>
Input graph edge
feature edge. relation between two entities: <relation name>
@@
Class node
prompt node. relation between two entities: < relation name >
Prompt node
prompt node. relation type prediction between the connected entities.
@@ B.1.7 Molecular Datasets
Node type
Text feature
Input graph node
feature node. atom. <element name>, <atom chirality>, degree of <atom degree>, formal charge of <formal charge>, num of hydrogen is <number of hydrogen>, num of radical electron is <number of radical electrons>, hybridization is <hybridization>, (is/is not) aromatic, (is/is not) in ring.
Input graph edge
feature edge. chemical bond. <bond type> bond, bond stereo is <bond stereo>, (is/is not) conjugated
@@
Class node
prompt node. molecule property description. <molecule property description. e.g. the molecule is effective to the following assay:…>
Prompt node
prompt node. Graph classification on molecule properties.
@@ B.2 Visualization of OFA dataset
In this section, we provide a visualization of all generated OFA datasets using the sentence transformer. Concretely, For each generated dataset, we randomly select 400 node embeddings and project it to 2 dimensions using TSNE. Note for molecular datasets, we include all node embeddings. The result is shown in Figure 4. We can see that the generated embeddings from different domains are successfully separated by the LLM as the embedding from molecular datasets, knowledge graphs, wiki pages, and citation networks are well separated in the visualization. Moreover, the LLM can even separate the citation network from different research domains. embeddings from Arxiv and Cora, which mainly contain papers from computer science are close to each other and far from the embeddings of Pubmed, which focus on biology. This result reveals one of the key rationales behind the success of the OFA. That is, by encoding the datasets from different domains into different sub-spaces, OFA allows one GNN to learn the information from different domains separately without impacting each other.
@@ E.2 Few-shot and Zero-shot Results
Results for the ogbn-arxiv and Cora can be found in Table 13 and Table 16, respectively. We can see that for the Cora dataset, OFA-joint-lr achieve better performance than OFA-ind-lr in all setting. This may indicate that the knowledge learned from other task levels like link- or graph-level can help the generalization of the model on Cora. For ogbn-arxiv dataset, the results for OFA-joint-lr and OFA-ind-lr are similar.

### v3

@@ B.1.2 PubMed
Cora is a citation network that contains papers and their citation relationship in the biomedical domain. The raw text data of the PubMed dataset was collected from the GitHub repository provided in Chen et al. (2023). All nodes and edges are similar to Cora dataset and we use the exact same processing procedure as Cora dataset. Because its original literature categories are diabetes, experimental/diabetes, type 1/diabetes, and type 2, which are overly simple and very difficult for our BERT-based LLM to distinguish, we asked ChatGPT to generate a detailed description of each category.
@@
Cora is a citation network that contains papers and their citation relationship collected from Arxiv platform. The raw text data of the ogbn-arxiv was collected using the same protocol as the GitHub repository provided in Prodigy (Huang et al., 2023). For ogbn-arxiv, we only evaluate the model on the node classification task. All nodes and edges are similar to Cora dataset and we use the exact same processing procedure as Cora dataset except that the description for each category is directly obtained from the Prodigy.
@@ B.1.4 Wiki-CS
Wiki-CS is an Internet link network with each node representing a Wikipedia page and each edge representing the reference link. The raw text of the Wiki-CS dataset was collected from the official website (Mernyei & Cangea, 2020). The raw text feature of a node is the name and content of an entry in Wikipedia. Each node’s label corresponds to the category of the entry. We evaluate Wiki-CS on node classification tasks. In table 9, we show a processed example on the Wiki-CS dataset using the processing protocol proposed in OFA.
@@
Node type Text feature
Input graph node feature node. wikipedia entry name: $<$entry name$>$. Entry content: $<$entry content$>$
Input graph edge feature edge. wikipedia page link.
@@
Class node prompt node. wikipedia entry category: $<$category name$>$
Prompt node prompt node. node classification of wikipedia entry category.
@@ B.1.5 FB15K237
FB15K237 is a knowledge graph that contains knowledge base relation triples and textual mentions of Freebase entity pairs. The raw text data of nodes in FB15K237 was collected from GitHub repository [^2] https://github.com/villmow/datasets_knowledge_embedding/tree/master. The raw text feature of a node is the name of the relation entity and its description. The raw text feature of an edge is the type of relation between two entities. We provide an example of processed data example in Table 10.
@@
Node type Text feature
Input graph node feature node. entity and entity description: <entity name>. <entity description>
Input graph edge feature edge. relation between two entities: <relation name>
@@
Class node prompt node. relation between two entities: < relation name >
Prompt node prompt node. relation type prediction between the connected entities.
@@ B.1.7 Molecular Datasets
Node type Text feature
Input graph node feature node. atom. <element name>, <atom chirality>, degree of <atom degree>, formal charge of <formal charge>, num of hydrogen is <number of hydrogen>, num of radical electron is <number of radical electrons>, hybridization is <hybridization>, (is/is not) aromatic, (is/is not) in ring.
Input graph edge feature edge. chemical bond. <bond type> bond, bond stereo is <bond stereo>, (is/is not) conjugated
@@
Class node prompt node. molecule property description. <molecule property description. e.g. the molecule is effective to the following assay:…>
Prompt node prompt node. Graph classification on molecule properties.
@@ B.2 Visualization of OFA dataset
In this section, we provide a visualization of all generated OFA datasets using the sentence transformer. Concretely, For each generated dataset, we randomly select 400 node embeddings and project it to 2 dimensions using TSNE. Note for molecular datasets, we include all node embeddings. The result is shown in Figure 4.
@@
We can see that the generated embeddings from different domains are successfully separated by the LLM as the embedding from molecular datasets, knowledge graphs, wiki pages, and citation networks are well separated in the visualization. Moreover, the LLM can even separate the citation network from different research domains. embeddings from Arxiv and Cora, which mainly contain papers from computer science are close to each other and far from the embeddings of Pubmed, which focus on biology. This result reveals one of the key rationales behind the success of the OFA. That is, by encoding the datasets from different domains into different sub-spaces, OFA allows one GNN to learn the information from different domains separately without impacting each other.
@@ D.1 Supervised learning experiments
For Cora-node, Pubmed-node, and ogbn-arxiv datasets, a different split is used. Specifically, for Cora-node and Pubmed-node, the split is obtained from Chen et al. (2023), where they split data into 10 folds and we use the first fold as our split. For ogbn-arxiv, in each experiment, we will randomly split data with a train/val/test ratio of 0.8/0.1/0.1. For Cora-node, Pubmed-node, ogbn-arxiv, WN18RR, and FB15K237, we rerun the baseline models 10 times and report the average performance. For other datasets, we report results from existing works.
@@
In few-shot graph construction, the text feature of class nodes differs significantly from that in supervised and zero-shot scenarios. Specifically, we omit category information and utilize a uniform text feature across all class nodes. This uniformity shifts the focus from learning class-specific information to enhancing the model’s ability to compare the query node with support nodes. Such comparisons are crucial for determining matches, which proves particularly effective in few-shot scenarios where the model must classify unseen classes.
@@ E.2 Few-shot and Zero-shot Results
Results for the ogbn-arxiv and Cora can be found in Table LABEL:tab:fs_arxiv_ablation and Table 16, respectively. We can see that for the Cora dataset, OFA-joint-lr achieve better performance than OFA-ind-lr in all setting. This may indicate that the knowledge learned from other task levels like link- or graph-level can help the generalization of the model on Cora. For ogbn-arxiv dataset, the results for OFA-joint-lr and OFA-ind-lr are similar.